OpenStudy (faman39):

Lydia received a parking ticket because she parked 12 feet from a fire hydrant. She had estimated her distance from the fire hydrant as being the minimum allowable distance, which is 15 feet. What was her percent error?

Directrix (directrix):

60% error - I worked this last week but will rework here.

OpenStudy (wasiqss):

(15-12)/15*100

OpenStudy (wasiqss):

answer is 20%

Directrix (directrix):

% error = | your result - accepted value | x 100 % accepted value ABS[ (16 - 10)/ 10 ] x 100 % = 6/10 x 100% = 60% error

OpenStudy (wasiqss):

see the values man!! its 12 and 15